OpenStudy (anonymous):

Identify intervals on which the function is increasing, decreasing, or constant. g(x) = 1 - (x - 6)2

OpenStudy (loser66):

calculate g'(x) , let it =0 to get the value of critical points. Plug each of them into g(x) to find value of g(x) at those points. Then, compare them, make table to know in which interval it increasing/decreasing /saddle.

OpenStudy (anonymous):

Is it -6 or +6 in the answer?

OpenStudy (loser66):

@cwrw238

OpenStudy (cwrw238):

can you find g'(x) or do you need help?

OpenStudy (anonymous):

Well, I know its going to be either increasing <-6 and decreasing > - 6 or increasing < 6 and decreasing > 6

OpenStudy (cwrw238):

g'(x) = -2(x - 6) = 0 at critical points x - 6 = 0 x = 6 to find if this is a maximum or a minimum we find g"(x) this = -2 negative gives a maximum so its is increasing for x < 6, maximum at x = 6 and decreasing for x > 6

OpenStudy (anonymous):

Thanks!

OpenStudy (cwrw238):

yw